OpenStudy (anonymous):

whats 4/28 in lowest terms

OpenStudy (diyadiya):

Both 4 & 28 are divisible by 4

OpenStudy (anonymous):

Hint: Find the prime factorizations of the numerator and denominator.\[\frac{4}{28}=\frac{2^2}{2^2\cdot 7}=?\]